If A =130 degrees and x= sin A +  cos A THEN


(A) X>0 (B) X<0 (C) X=0 (D) X<=0


eXPLAIN PLS...ANS GIVEN IS (A)...

Grade:12th Pass

2 Answers

omi more
11 years ago

x= sin(130) +cos(130)= sin 70* + (-cos 70*)    =>x>0 (since sin 70* would be greater than cos 70*)

Venkat
4 years ago
X= Sin130 + Cos130 =Sin(180-50) + Cos(180-50)
                                     =Sin50 - Cos50
                                     =Sin50 - Cos(90-40)
                                     =Sin50 - Sin40
                                     =Sin(45+5) - Sin(45-5)
                                     =2Cos45.Sin5 
                    Since, the value of 2cos45.sin5 is positive 
                         the value of X is also positive.